# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v3

@@ Abstract
Accurate forecasting of epidemic infection trajectories is crucial for safeguarding public health. However, limited data availability during emerging outbreaks and the complex interaction between environmental factors and disease dynamics present significant challenges for effective forecasting. In response, we introduce CAPE, a novel epidemic pre-training framework designed to harness extensive disease datasets from diverse regions and integrate environmental factors directly into the modeling process for more informed decision-making on downstream diseases. Based on a covariate adjustment framework, CAPE utilizes pre-training combined with hierarchical environment contrasting to identify universal patterns across diseases while estimating latent environmental influences. We have compiled a diverse collection of epidemic time series datasets and validated the effectiveness of CAPE under various evaluation scenarios, including full-shot, few-shot, zero-shot, cross-location, and cross-disease settings, where it outperforms the leading baseline by an average of 9.9% in full-shot and 14.3% in zero-shot settings. The code will be released upon acceptance.
Machine Learning, ICML
@@ 1 Introduction
Infectious disease outbreaks consistently challenge public health systems, affecting both individual well-being and economic stability (Nicola et al., 2020). Effective management of these outbreaks hinges on accurate epidemic forecasting, which involves predicting future incidences like infection cases and hospitalizations (Liu et al., 2024b; Wan et al., 2024; Adhikari et al., 2019). Over the years, various models have been developed to address this need. These include mechanistic models like SIR (Cooper et al., 2020) and statistical models like ARIMA (Sahai et al., 2020; Kontopoulou et al., 2023), as well as advanced machine learning methods such as LSTM and GRU (Shahid et al., 2020), which have proven instrumental in forecasting disease spread and supporting informed public health decision-making.
Despite the advancements, current models are typically trained for specific diseases within particular geographic regions, limiting their ability to integrate insights from diverse sources spanning multiple pathogens and spatial contexts. This narrow focus can impede a comprehensive understanding of disease dynamics and the design of effective outbreak responses, especially during novel or emergent outbreaks when observations are typically scarce. Taking Influenza, Measles, and Dengue as examples (see dataset details in Appendix A.4), the observed performance gap in Figure 1 between a simple ARIMA model and more complex deep learning models, especially Transformer-based models, highlights a key limitation: more complex models often struggle in epidemic forecasting when prior knowledge is unavailable and observational data is scarce. Nevertheless, given the extensive and diverse outbreak data collected over decades and across various geographies, pre-training on such a broad dataset could potentially enable the development of more generalizable models with greater applicability and adaptability across different pathogens and contexts. This raises an important question: Can we leverage lessons from diverse historical disease time series to enhance epidemic forecasting accuracy for more complex transformer-based models?
To address the above question, we draw inspiration from the success of large pre-trained transformer-based models (Zhao et al., 2023) and develop a pre-trained epidemic forecasting model using extensive disease time series data to distill generalizable knowledge across pathogens and contexts. The pre-trained model can be subsequently fine-tuned for specific diseases or geographical regions. While it is possible to adapt general time series foundation models (Liang et al., 2024; Ma et al., 2024) to epidemic forecasting, their pre-trained corpus mostly consists of non-epidemic data, which may not accurately capture epidemic dynamics and infection trajectories, potentially degrading forecasting accuracy.
Furthermore, although an early effort has been made in epidemic pre-training (Kamarthi & Prakash, 2023), it overlooks critical external factors, which we refer to as environment throughout this paper. These factors such as temperature, humidity, public health policies, and interventions are well-known to influence the dynamics of disease spread in space and time (Lau et al., 2020b; Chen & Hsieh, 2012). As the environment influences both historical infection trajectories and future disease spread, it can be seen as an important confounder from the perspective of causal inference (Zhou et al., 2023; Jiao et al., 2024), which hinders the learning of true causal relations between the historical infections and future outbreaks, as shown in the structured causal model in Figure 2. Driven by the changes in the environment, a major obstacle arises as the shift in the distribution of infection trajectories between training and test datasets  (Wu et al., 2024). Insufficient consideration of such distribution shifts can obscure the relationship between historical infection data and future predictions (for a detailed discussion, see Appendix A.15), compromising a model’s ability to make accurate forecasts. As such, it is crucial to disentangle the influence of changing environments from other more intrinsic factors (e.g., a pathogen’s infection rate) affecting disease transmission dynamics. Yet, exact and explicit mechanisms by which the environment influences the disease dynamics of a particular pathogen are often not fully understood or observed, which necessitates an implicit modeling approach to identify and separate these latent environmental influences.
Figure 1: Performance gap between the ARIMA model and various non-pretrained deep learning models across selected datasets and horizons. Positive values indicate that the deep learning models underperform relative to ARIMA, highlighting the limitations of complex architectures without pretraining or prior knowledge.
Figure 2: Structural causal model for epidemic forecasting, where $Z$ refers to the environment as a confounder. $X_{sp}$ and $X_{ca}$ are the spurious and causal factors of the input.
Our Solution. We propose Covariate-Adjusted Pretraining for Epidemic forecasting (CAPE) to capture universal, environment-invariant disease patterns from historical data (Figure 3). CAPE tackles data scarcity and distribution shift by integrating pre-training with implicit environment modeling, grounded in causal analysis. Our contributions include:
  1. As one of the first epidemic pre-trained models, we propose a generalizable pre-training framework that adjusts for covariates in epidemic time series data, disentangling intrinsic disease dynamics from environmental influences and yielding environment-invariant representations.
  2. We create a comprehensive testbed for evaluating pre-trained epidemic models by assembling epidemic time series datasets across various diseases and regions. This enables thorough assessment across various scenarios, including few-shot, zero-shot, cross-location, and cross-disease settings.
  3. We demonstrate the potential of epidemic pre-training, showcasing superior performance across downstream datasets and settings. Notably, CAPE surpasses the best baseline by an average of 9.9% in the full-shot setting and 18.1% in the zero-shot setting across all tested downstream datasets.
  4. We provide an in-depth analysis of how pre-training and environment modeling affect downstream performance and mitigate the impact of distribution shifts.
@@ 2 Related Work and Problem Definition
Epidemic Forecasting Models. Traditionally, epidemic forecasting employs models like ARIMA (Sahai et al., 2020), SEIR (He et al., 2020), and VAR (Shang et al., 2021). ARIMA predicts infections by analyzing past data and errors, SEIR models population transitions using differential equations, and VAR captures linear inter-dependencies by modeling each variable based on past values. Recently, deep learning models—categorized into RNN-based, MLP-based, and transformer-based—have surpassed these methods. RNN-based models like LSTM (Wang et al., 2020) and GRU (Natarajan et al., 2023) use gating mechanisms to manage information flow. MLP-based models use linear layers (Zeng et al., 2023) or multi-layer perceptrons (Borghi et al., 2021; Madden et al., 2024) for efficient data-to-prediction mapping. Transformer-based models (Wu et al., 2021; Zhou et al., 2021, 2022) apply self-attention to encode time series and generate predictions via a decoder. However, these models are limited as they typically utilize data from only one type of disease without considering valuable insights from diverse disease datasets.
@@
Pre-trained Time Series Models. To enable few-shot or zero-shot capabilities, transformer-based models often employ pre-training on large datasets, which typically use masked data reconstruction (Zerveas et al., 2021; Rasul et al., 2023) or promote alignment across different contexts (Fraikin et al., 2023; Zhang et al., 2022; Yue et al., 2022). For example, PatchTST (Nie et al., 2022) segments time series into patches, masks some, and reconstructs the masked segments. Larger foundational models like MOMENT (Goswami et al., 2024) aim to excel in multiple tasks (e.g., forecasting, imputation, classification) but require substantial data and computational resources. In epidemic contexts, Kamarthi et al. (Kamarthi & Prakash, 2023) pre-train a model on various diseases, improving downstream performance and highlighting pre-training’s potential in epidemic forecasting. However, these works overlook environmental influence and zero-shot ability in epidemic forecasting and lack a deep analysis of how pre-training materials impact downstream performance. In this study, we introduce environment modeling and conduct a thorough analysis of these questions (see A.5 for more discussions).
Problem Definition. Given a historical time series input: $\mathbf{x}\in\mathbb{R}^{T\times 1}$, where $T$ is the size of lookback window, the goal of epidemic forecasting is to map $\mathbf{x}$ into target trajectories (e.g. infection rates): $\mathbf{y}\in\mathbb{R}^{h}$, where $h$ denotes the size of the forecast horizon. We define $X$ and $Y$ as the random variables of input and target, respectively. During pre-training, a representation function $g_{\theta}:\mathbb{R}^{T\times 1}\to\mathbb{R}^{T\times d}$, where $d$ denotes the dimension of the latent space and $\theta$ being the parameter of the model, extracts universal properties from a large collection of epidemic time series datasets $\mathcal{D}_{\text{pre}}=\{D_{1}^{\prime},D_{2}^{\prime},\dots,D_{S}^{\prime}\}$. Then, a set of self-supervised tasks $\mathcal{T}_{\text{pre}}=\{\mathcal{T}_{i}\}_{i=1}^{R}$ is defined, and each $\mathcal{T}_{i}$ transforms a sample $\mathbf{x}\sim\mathcal{D}_{\text{pre}}$ into a new input-label pair: $(\mathbf{\tilde{x}},\mathbf{\tilde{y}})$, and optimizes a loss $\mathcal{L}_{\mathcal{T}_{i}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{pre}}}[\ell_{\mathcal{T}_{i}}(h_{\psi}(g_{\theta}(\mathbf{\tilde{x}})),\mathbf{\tilde{y}})]$, with $\ell_{\mathcal{T}_{i}}$ being the task-specific metric and $h_{\psi}$ the task-specific head.
@@ 3 Proposed Method
Pre-training models for epidemic forecasting using data from diverse contexts presents two primary challenges. First, the inherent heterogeneity stemming from varied environmental conditions (e.g., public health interventions, seasonal factors, population behaviors) necessitates a model architecture capable of accounting for these influences. Second, the pre-training objectives must be designed to foster the learning of robust representations that capture these intrinsic dynamics and generalize effectively across diverse epidemiological scenarios, including novel or data-scarce situations. We will elaborate on our solutions in the following subsections to address these challenges.
3.1 Causal Analysis for Epidemic Forecasting
In epidemic forecasting, environmental factors, such as seasonal changes and population mobility, are known to confound the relationship between historical infection patterns and future disease spread. To address this critical challenge in understanding true transmission dynamics, we draw inspiration from causal inference (Zhou et al., 2023; Jiao et al., 2024) and introduce a Structural Causal Model (Figure 2) where we treat the environment $Z$ as an unobserved confounder that influences both the independent variable (e.g., historical infections $X$) and the dependent variable (e.g., future infections $Y$). Furthermore, based on the causal decomposition assumption (Mao et al., 2022), we know $X$ is composed of (1) a spurious factor $X_{sp}$ that is environment-dependent, and (2) a causal factor $X_{ca}$ representing the underlying, more intrinsic disease transmission dynamics (e.g., pathogen infectivity) that remains environment-independent. While $X$ as a whole is observed, only $X_{ca}$ has a causal relation with target $Y$. Here, we assume that a finite set of critical factors drives the epidemic dynamics:
Assumption 3.1. The environment variable $Z$ follows a categorical distribution $p(Z)$ and takes on one of $K$ discrete environmental states, denoted as $z_{k}$. Each state $z_{k}$ is associated with a unique latent representation $\mathbf{e}_{k}\in\mathbb{R}^{h_{e}}$, capturing the unique features specific to that environment.
In constructing a predictive model for input $\mathbf{x}$, we define $\hat{Y}$ as the predicted time series $\hat{\mathbf{y}}$ and model the predictive distribution $p_{\Theta}(\hat{Y}|X)$ using $f_{\Theta}(\mathbf{x})=h_{\psi}(g_{\theta}(\mathbf{x}))$, where $\Theta=\{\theta,\psi\}$. Training typically involves maximizing the log-likelihood of $p_{\Theta}(\hat{Y}|X)$, which in practice translates to minimizing the errors over the pre-training dataset $\mathcal{D}_{\text{pre}}$: $\Theta^{*}=\arg\min_{\Theta}-\frac{1}{|\mathcal{D}_{\text{pre}}|}\sum\nolimits_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{\text{pre}}}\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}$. As the environment $Z$ impacts the distribution of the observed data through $p(X,Y|Z)=p(X|Z)p(Y|X,Z)$, we formulate the following objective:
$\begin{split}\Theta^{*}=&\arg\min\nolimits_{\Theta}-\mathbb{E}_{p(Z)}[\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p(Y,X|Z)}[\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}]].\end{split}$ (1)
The above equation suggests that the optimal $\Theta^{*}$ depends on the environment distribution $p(Z)$. If we simply maximize the likelihood $p_{\Theta}(\hat{Y}|X)$, the confounding effect of $Z$ on $X$ and $Y$ will mislead the model to capture the shortcut predictive relation between the input and the target trajectories, which necessitates implicit modeling of the environment during pre-training. Given that input infection trajectories inherently reflect the influence of the environment, it is crucial to develop mechanisms that disentangle the correlations between infection trajectories and environmental factors.

### v4

# Pre-training Epidemic Time Series Forecasters with Compartmental Prototypes

@@ Abstract
Accurate epidemic forecasting is crucial for outbreak preparedness, but existing data-driven models are often brittle. Typically trained on a single pathogen, they struggle with data scarcity during new outbreaks and fail under distribution shifts caused by viral evolution or interventions. However, decades of surveillance data from diverse diseases offer an untapped source of transferable knowledge. To leverage the collective lessons from history, we propose CAPE, the first open-source pre-trained model for epidemic forecasting. Unlike existing time series foundation models that overlook epidemiological challenges, CAPE models epidemic dynamics as mixtures of latent population states, termed compartmental prototypes. It discovers a flexible dictionary of compartment prototypes directly from surveillance data, enabling each outbreak to be expressed as a time-varying mixture that links observed infections to latent population states. To promote robust generalization, CAPE combines self-supervised pre-training objectives with lightweight epidemic-aware regularizers that align the learned prototypes with epidemiological semantics. On a comprehensive benchmark spanning 17 diseases and 50+ regions, CAPE significantly outperforms strong baselines in zero-shot, few-shot, and full-shot forecasting. This work represents a principled step toward pre-trained epidemic models that are both transferable and epidemiologically grounded.
@@ 1 Introduction
Infectious disease outbreaks pose a persistent threat to global public health and economic stability (Nicola et al., 2020). Effective outbreak management relies on accurate epidemic forecasting—the prediction of future cases, hospitalizations, and other critical metrics (Liu et al., 2024b; Wan et al., 2024; Adhikari et al., 2019). A wide range of models have been developed to provide these crucial forecasts, which generally fall into two categories. Mechanistic models, such as the classic Susceptible-Infected-Recovered (SIR) (Cooper et al., 2020) approach, are grounded in epidemiological principles; they divide a population into compartments that represent distinct population states (e.g., susceptible, infectious, recovered) and use differential equations to explicitly model flows between these states. In contrast, modern machine learning methods like LSTMs (Shahid et al., 2020) learn complex patterns directly from historical data, offering greater flexibility without imposing a predefined structure.
However, these data-driven forecasters are often trained for a single pathogen in a specific region. This narrow scope makes them brittle: they face acute data scarcity during the critical early stages of a novel outbreak, and they fail under distribution shifts induced by viral evolution, behavioral change, or policy interventions. At the same time, decades of surveillance across diverse pathogens and geographies remain an untapped source of transferable structure. Motivated by the success of large pre-trained models in language, vision, and time-series domains (Zhao et al., 2023), we ask: Can we build a large pre-trained epidemic forecaster that learns from the collective history of infectious diseases to improve generalization and robustness?
Simply applying a general time series foundation model (Liang et al., 2024) is insufficient, as it overlooks core epidemiological challenges: (1) Structural heterogeneity: Pathogens follow different effective compartmental progressions (e.g., SIR vs. SEIR (He et al., 2020)), so a single fixed mechanism cannot transfer broadly across diseases and regions. (2) Hidden population states: Surveillance data records only reported infections, while important states such as exposure, susceptibility, and immunity are not directly observed. (3) Distribution Shifts: Interventions, behavioral changes, and pathogen evolution induce abrupt non-stationarities, often when outbreak histories are shortest. These properties demand powerful epidemic pre-trained models that can adapt to diverse pathogens, disentangle hidden population states, and remain robust under shifts.
Our Solution. We introduce CAPE (CompArtment Pre-training for Epidemics), a pre-trained framework that learns epidemic dynamics as a mixture of latent population states, termed compartmental prototypes. (1) To address structural heterogeneity, rather than relying on a rigid, pre-defined compartmental structure, CAPE discovers a flexible dictionary of latent compartments directly from data. Each outbreak sequence is modeled as a mixture that varies in time in these prototypes, linking observed infections to latent population states. (2) To handle hidden drivers, the learned mixtures act as proxies for unobserved states such as susceptibility, disentangling latent population dynamics from noisy observed case counts. (3) To address distribution shifts, CAPE employs two self-supervised pre-training strategies to encourage robust representations that generalize under non-stationarities and scarce data. In addition, we further propose lightweight epidemic-aware regularizers to align learned prototypes with epidemiological semantics. Our contributions include:
  1. (1)
    Pre-training framework for epidemic time series forecasting: We introduce the first open-source pre-training framework  [^1] https://anonymous.4open.science/r/CAPE_ICLR26-A041/ for epidemic forecasting. It learns latent compartmental prototypes directly from time series, guided by several epidemic-aware losses that regularize the model’s predictions and learned prototype representations.
  2. (2)
    Comprehensive benchmark for epidemic pre-training: We assemble a diverse pre-training and evaluation suite, spanning 17 diseases across 50+ regions for pre-training and 5 downstream datasets covering 4 challenging settings (zero-shot, few-shot, cross-location, and cross-disease).
  3. (3)
    State-of-the-art performance in diverse forecasting settings: We demonstrate the effectiveness of our pre-trained model, which significantly outperforms existing benchmarks by an average of 6.3% lower average MSE in the full-shot setting and 10.3% lower average MSE in the few-shot setting across all tested downstream datasets.
  4. (4)
    In-depth analysis: We conduct extensive analyses to provide insights into how the learned latent prototypes improve forecasting accuracy and show that pre-training effectively learns the representation of diverse diseases and mitigates the impact of distribution shifts.
@@ 2 Related Work and Problem Definition
Epidemic Forecasting Models. Traditionally, epidemic forecasting employs models like ARIMA (Sahai et al., 2020), SEIR (He et al., 2020), and VAR (Shang et al., 2021). ARIMA predicts infections by analyzing past data and errors, SEIR models population transitions using differential equations, and VAR captures linear inter-dependencies by modeling each variable based on past values. Recently, deep learning models—categorized into RNN-based, MLP-based, and transformer-based—have surpassed these methods. RNN-based models like LSTM (Wang et al., 2020) and GRU (Natarajan et al., 2023) use gating mechanisms to manage information flow. MLP-based models use linear layers (Zeng et al., 2023) or multi-layer perceptrons (Borghi et al., 2021; Madden et al., 2024) for efficient data-to-prediction mapping. Transformer-based models (Wu et al., 2021; Zhou et al., 2021; 2022) apply self-attention to encode time series and generate predictions via a decoder. However, these models are limited as they typically utilize data from only one type of disease without considering valuable insights from diverse disease datasets.
@@
Pre-trained Time Series Models. To enable few-shot or zero-shot capabilities, transformer-based models often employ pre-training on large datasets, which typically use masked data reconstruction (Zerveas et al., 2021; Rasul et al., 2023) or promote alignment across different contexts (Fraikin et al., 2023; Zhang et al., 2022; Yue et al., 2022). For example, PatchTST (Nie et al., 2022) segments time series into patches, masks some, and reconstructs the masked segments. Larger foundational models like MOMENT (Goswami et al., 2024) aim to excel in multiple tasks (e.g., forecasting, imputation, classification) but require substantial data and computational resources. In epidemic contexts, Kamarthi et al. (Kamarthi & Prakash, 2023) pre-train a model on various diseases, improving downstream performance and highlighting pre-training’s potential in epidemic forecasting. However, the complete implementation is not publicly available. Moreover, existing approaches overlook hidden compartmental influence and zero-shot ability in epidemic forecasting and lack a deep analysis of how pre-training materials impact downstream performance. In this study, we introduce latent compartment modeling and conduct a thorough analysis of these questions (see A.6 for more discussions).
[…]
@@ 3 Proposed Method
Our pre-training framework is designed to overcome the core challenges of structural heterogeneity, hidden drivers, and distribution shifts inherent in epidemic forecasting. We address these issues through two main contributions: (1) a flexible model architecture that learns latent compartmental prototypes directly from observational data, and (2) a set of epidemic-aware pre-training objectives that guide the model to learn robust, generalizable representations. We will elaborate on these architectural and objective-based solutions in the following subsections.
Figure 1: (a) CAPE encoder and environment estimator with latent representations; (b) Hierarchical environment contrasting for temporal and environment representations; (c) Random masking and reconstruction with environment estimation to capture universal patterns; (d) EM algorithm to iteratively optimize model parameters and environment representations. (e) Epidemic-aware regularization, including losses for monotonic and non-monotonic dynamics.
3.1 Modeling latent compartmental prototypes
Temporal Backbone. Following the prior work on patch-based time series modeling  (Nie et al., 2022), we segment the input sequence of infection counts $\mathbf{x}$ into non-overlapping temporal patches, $\mathbf{x}=[\mathbf{x}_{1},\dots,\mathbf{x}_{C}]$, where each patch $\mathbf{x}_{c}\in\mathbb{R}^{T/C}$. This patching strategy enables the model to capture local temporal patterns. A standard self-attention encoder, $f_{\text{enc}}$, then processes these patches to learn long-range temporal dependencies, producing a contextualized representation $\mathbf{h}^{(l)}_{c}=f_{enc}(\mathbf{x}^{l}_{c})$ for each patch $c$ at a given layer $l$.
Compartmental Prototypes Learning. Different diseases exhibit different progression patterns: some show a simple rise-and-fall in cases, while others involve additional hidden stages such as incubation periods. In epidemiological terms, this corresponds to differences in compartmental structures (e.g., SIR (Cooper et al., 2020) vs. SEIR (He et al., 2020)). Epidemic forecasting is therefore challenged by structural heterogeneity and hidden population states: different structures generate diverse dynamics, while only a subset of compartments are directly observable. Classic compartmental models impose a rigid, pre-defined structure that cannot adapt to various types of diseases and scenarios. To address this, we move beyond fixed models and propose a framework that learns to represent epidemic dynamics as a dynamic mixture of latent population states, which we term compartmental prototypes. Our approach is analogous to learning a vocabulary of core epidemiological behaviors directly from data. Each outbreak is expressed as a time-varying mixture over these prototypes. For example, during the growth phase of influenza, the model may infer that roughly 30% observed case increases are explained by infectious-like prototypes, while the rest reflects susceptible depletion and recovery. Formally, we initialize a set of $K$ learnable embeddings, $\mathbf{E}=\{\mathbf{e}_{k}\}_{k=1}^{K}\in\mathbb{R}^{K\times d}$, where each $\mathbf{e}_{k}$ is a ”prototype” representing a basic compartment. For any given time window (a patch $c$), the model’s crucial task is to determine the contributions of these prototypes in representing the current epidemic time series. We accomplishes this by inferring a set of mixture weights, $\bm{\pi}_{c}=[\pi_{1,c},\dots,\pi_{K,c}]$, using a cross-attention mechanism between the patch’s representation $\mathbf{h}_{c}^{(l)}$ and the full set of prototype embeddings $\mathbf{E}$:
where $\mathbf{W}_{k}^{(l)}$ and $\mathbf{W}_{s}^{(l)}$ are learnable linear projections. $\bm{\pi}_{c}$ quantifies the contribution of each compartmental prototype in representing the current patch, forming a regularized and robust representation for forecasting. The patch representation is then updated by taking a weighted sum over the Hadamard product ($\odot$) of the patch representation and the compartment embeddings. This allows the model to modulate the observed time series data with the inferred underlying dynamics. The layer-wise update is defined as:
$\mathbf{x}^{(l+1)}_{c}=\sigma\left(\mathbf{W}_{f}^{(l)}\sum\nolimits_{k=1}^{K}\pi_{k,c}^{(l)}\left[f_{enc}({\mathbf{x}_{c}^{(l)}})\odot\mathbf{e}_{k}\right]\right),$ (2)
where $\sigma$ represents a feed-forward block containing the projection $\mathbf{W}_{f}^{(l)}$. After stacking $L$ such layers, a final task-specific linear head, $h_{\psi}$, maps the resulting representations $\mathbf{x}^{(L)}$ to the target prediction $\mathbf{\hat{y}}=h_{\psi}(\mathbf{x}^{(L)})$.
3.2 Self-supervised Pre-training
To learn robust representations that can withstand the distribution shifts and data scarcity common in epidemics, we employ two self-supervised pre-training objectives designed to capture universal patterns across diverse time series.
Masked Time-Series Reconstruction. We use a masked autoencoding task to teach the model the underlying grammar of epidemic curves. By randomly masking a fraction (e.g., 30%) of the input patches and training the model to reconstruct the original series, we force it to learn meaningful temporal interpolations. The objective is to minimize the Mean Squared Error, $\mathcal{L}_{\text{recon}}=\text{MSE}(\hat{\mathbf{x}},\mathbf{x})$. This builds resilience to the noisy and incomplete data often encountered during chaotic outbreak periods, improving the model’s fundamental forecasting capabilities.
Contrastive Learning for Compartmental Prototypes. A key challenge during distribution shifts is that epidemic curves can become highly non-stationary, and superficially similar patterns might arise from vastly different underlying dynamics. To prevent our model from learning spurious correlations, we introduce a contrastive objective that regularizes the compartmental prototype mechanism itself. The goal is to ensure that the inferred contributions of compartments ($\bm{\pi}_{c}$) are both consistent and discriminative. Specifically, we enforce two conditions: (1) two different augmented views of the same time-series patch should be mapped to a similar mixture of compartmental prototypes (positive pairs), and (2) patches from different, epidemiologically distinct contexts should be mapped to dissimilar compartmental prototypes (negative pairs). This pushes the model to focus on the essential, underlying dynamics captured by the prototypes, rather than overfitting to superficial noise. The patch-wise contrastive loss is defined as:
$\mathcal{L}_{\text{CL}}(j,c)=$ $-\mathbf{X}_{(j,c)}\cdot\mathbf{X}^{\prime}_{(j,c)}+\log\left(\sum\nolimits_{b\in B}\exp\left(\mathbf{X}_{(j,c)}\cdot\mathbf{X}^{\prime}_{(b,c)}\right)+\mathbb{I}_{j\neq b}\exp\left(\mathbf{X}_{(j,c)}\cdot\mathbf{X}_{(b,c)}\right)\right)$ (3)
$+\log\left(\sum\nolimits_{t\in\Omega}\exp\left(\mathbf{X}_{(j,c)}\cdot\mathbf{X}^{\prime}_{(j,t)}\right)+\mathbb{I}_{c\neq t}\exp\left(\mathbf{X}_{(j,c)}\cdot\mathbf{X}_{(j,t)}\right)\right),$
where $B$ is the batch, $\Omega$ is the set of overlapping patches, and $\mathbb{I}$ is the indicator function.
3.3 Epidemic-aware Regularization
To ensure our compartmental prototypes learn epidemiologically plausible dynamics, we introduce three regularization terms that instill prior knowledge from classic mechanistic models. These regularizers help disentangle the learned prototypes, encouraging them to represent distinct, interpretable dynamics (e.g., monotonic vs. non-monotonic). We apply these regularizers with a small weight during pre-training to gently guide representation learning, and with a larger weight during fine-tuning to specialize the model to a specific pathogen.
Monotonic Dynamics. Certain compartments, such as Susceptible or Recovered, typically exhibit monotonic behavior (Nguyen et al., 2023). To enforce this, we introduce a monotonic loss. For a compartment $k$ with an expected monotonic trend, let $\bm{\pi}_{k}=[\pi_{k,1}^{(L)},\dots,\pi_{k,C}^{(L)}]$ be its mixture weights sequence from the final layer. The monotonic decreasing loss is:
$\mathcal{L}_{\text{mono}}=\frac{1}{C-1}\sum_{c=2}^{C}\text{ReLU}(\pi_{k,c}^{(L)}-\pi_{k,c-1}^{(L)}+\epsilon),$ (4)
where $\epsilon>0$ is a small tolerance. The ReLU function penalizes only violations of the expected trend. For an increasing trend, the terms in the parentheses are swapped. In practice, we constrain two prototypes with increasing and decreasing monotonic penalties.
Non-monotonic Dynamics. While monotonic constraints are simple and effective for some prototypes, the dynamics of active infections are complex and non-monotonic. Therefore, simply applying a specific predefined pattern to such prototypes can be ineffective. Instead, we regulate their behavior using one of the most fundamental principles in epidemiology: the basic reproduction number, $R_{0}$, which quantifies a pathogen’s intrinsic transmissibility. Our goal is to ensure that the infectious dynamics learned by our model correspond to a plausible $R_{0}$ for the disease being modeled. To achieve this, we introduce a method to compute a differentiable proxy for $R_{0}$ directly from our model’s learned representations. We adapt the classic Next Generation Matrix (NGM) method (Diekmann et al., 2010), denoted as $\mathbf{G}$, based on the Disease-Free Equilibrium time series input. Then, the $R_{0}$ is defined as the spectral radius of $\mathbf{G}$: $\hat{R}_{0}^{\text{raw}}=\max_{j}\|\lambda_{j}(\mathbf{G})\|$, which has the following lower- and upper-bounds (see proof in Appendix A.4.2):
$\frac{\sigma_{\text{min}}(\mathbf{F})}{\sigma_{\text{max}}(\mathbf{V})}\leq\max_{j}\|\lambda_{j}(\mathbf{G})\|=\max_{j}\|\lambda_{j}(\mathbf{F}\mathbf{V}^{-1})\|\leq\frac{\sigma_{\text{max}}(\mathbf{F})}{\sigma_{\text{min}}(\mathbf{V})},$ (5)
where $\mathbf{F}$ is the Jacobian of the rates of flows from uninfected to infected classes evaluated at the disease-free equilibrium, and $\mathbf{V}$ is the Jacobian of the rates of all other flows to and from infected compartments, $\lambda_{j}(\mathbf{G})$ are the eigenvalues of $\mathbf{G}$ and $\sigma_{\text{min}}(\mathbf{V})$ is the smallest singular value of $\mathbf{V}$. Since computing the inverse of matrix $\mathbf{V}$ is not always numerically stable, we approximate the lower- and upper-bound of spectral radius via the singular value ratios $\frac{\sigma_{\text{min}}(\mathbf{F})}{\sigma_{\text{max}}(\mathbf{V})}$ and $\frac{\sigma_{\text{max}}(\mathbf{F})}{\sigma_{\text{min}}(\mathbf{V})}$. Further details are provided in Appendix A.4.1, and here we provide a pseudo code for calculating $R_{0}$ in Table 1.
Algorithm 1 NGM-Proxy($R_{0}$): Differentiable $R_{0}$ bounds
1:Encoder $f_{\mathrm{enc}}$, prototypes $\mathbf{E}$, estimator $g$, mix operator $\phi$, disease $d$, range $R^{\mathrm{lo}}_{0},R^{\mathrm{hi}}_{0}]$.
2:Estimates $(\widehat{R}^{\mathrm{lo}}_{0},\widehat{R}^{\mathrm{hi}}_{0})$, loss $\mathcal{L}_{R_{0}}$.
3:DFE: Compute DFE embedding $\mathbf{E}_{\mathrm{DFE}}\!\leftarrow\!f_{\mathrm{enc}}(\mathbf{X}_{\mathrm{DFE}})$ and weights $\bm{\pi}^{*}\!\leftarrow\!\text{softmax}(\mathbf{E}_{\mathrm{DFE}}\mathbf{E}^{\top})$.